OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x^2 + 2x + 5 =0

OpenStudy (anonymous):

a-1, b=2, c=5 D=2^2-4(1)(5) D=4-20=-16<0. So Roots Imaginary